OpenStudy (anonymous):

Find the simplified form of the expression. Give your answer in scientific notation. (8 x 10^7) (7 x 10^4) a. 1.5 × 10^12 b. 5.6 × 10^12 c. 1.5 x 10^29 d. 5.6 x 10^29

OpenStudy (anonymous):

You multiply the coefficients, and add the explonents

OpenStudy (anonymous):

so I multiply (8x10) and (7x10) and then when I get done with that ad the7 and 4?

OpenStudy (anonymous):

Yeah, what do you get?

OpenStudy (anonymous):

You can just multiply 8 and 7.

OpenStudy (anonymous):

56

OpenStudy (anonymous):

And alltogether?

OpenStudy (anonymous):

sorry I took so long to reply my computer was acting crazy....so I add the exponents now

OpenStudy (anonymous):

I think it's the website going crazy.

OpenStudy (anonymous):

Did you get 56x10^11?

OpenStudy (anonymous):

I don't know but yeah I was about to type that

OpenStudy (anonymous):

Now you just make it a scientific notation by moving the decimal point, but note that you also have to increase/decrease the exponent depends on where you move it.

OpenStudy (anonymous):

560x10^10 = 56x10^11 = 5.6x10^12 = 0.56x10^13

OpenStudy (anonymous):

but that is not one of my answer choices im not sure what to do.

OpenStudy (anonymous):

Not sure if that's a delayed message, but look above.

OpenStudy (anonymous):

Or below? I'm not sure what is going on with the website.

OpenStudy (anonymous):

so its 5.6x10^12 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you so much